# Reconfigurable dissipative entanglement between many spin ensembles: from robust quantum sensing to many-body state engineering

### v1

@@ III.3 Impact of single-spin dissipation during state preparation
In the presence of such unwanted dissipation, the relaxation timescale $t_{\rm ss}$ for the ideal dynamics to reach the steady state $|\psi_{\rm ss}^{(2)}\rangle$ must not be too large, otherwise the steady-state spin squeezing will be strongly degraded once we introduce $\gamma$. We define $t_{\rm ss}$ such that the ideal ($\gamma=0$) evolution yields a high fidelity with the steady state for $t>t_{\rm ss}$. More specifically  $F(t)=\mathrm{tr}\Big(|\psi_{\rm ss}^{(2)}\rangle\langle\psi_{\rm ss}^{(2)}|\,\hat{\rho}(t)\Big)$ satisfies $F(t)>1-\epsilon$ for $t>t_{\rm ss}$, where $\epsilon$ is a small positive number (we choose $\epsilon=10^{-3}$). Here, $\hat{\rho}(0)$ can be any specific state easy to initialize in an experiment. One finds that there is a fundamental tradeoff between the speed of the dynamics and the amount of steady-state squeezing and entanglement. Similar to Refs. 45, 46, one can quantify this with a lower bound on the relaxation time (see App. B),
@@
In Fig. 3(b), we compute $t_{\rm ss}$ numerically using the quantum-jump method [47] and compare with $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}$. We use an initial state $|J=2S,m=-2S\rangle$, i.e., all spins pointing down, and then minimize $t_{\rm ss}$ by varying the ratio $\Omega/\Gamma$ while keeping $\Gamma$ and $\Delta/\Omega$ fixed. Note that optimal $t_{\rm ss}$ can be achieved when $\Omega\gtrsim S\Gamma$. In this case, we find $\Gamma t_{\rm ss}\approx 1/\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}$, which is a constant factor away from the lower bound (see Eq. (13)). As we discussed in Sec. III.2, $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}\sim S\Delta/\Omega$, we thus have $t_{\rm ss}\sim\Omega/(S\Gamma\Delta)$.
@@ III.4 Robustness against common-mode noise during interrogation
Figure 4: (a) Schematic for differential sensing in the presence of common phase noise. We first apply the quantum channel $\mathcal{E}_{12}$ (see Eq. (3)) to stabilize the steady state $|\psi_{\rm ss}^{(2)}\rangle$. The system then evolves under the differential phase $\phi$. We next apply a global $\pi/2$ pulse (random common phase $\theta$ due to laser noise) and then perform projective measurement of each ensemble in the $z$ basis to obtain the excitation fractions $p_{1}$ and $p_{2}$. (b) Color-scale plot of the probability distribution for measurement outcomes $(p_{1},p_{2})$, for $S=50$, $\Delta/\Omega=1/\sqrt{S}$ and $\phi=0.1$. The points $(p_{1},p_{2})$ with peak probability lie on an ellipse whose form depends on $\phi$. (c) Classical Fisher information $F_{\phi}$ for ellipse fitting (solid lines). The dashed lines are the corresponding quantum Fisher information $\mathcal{F}_{\phi}$. $F_{\phi}$ is roughly approaching $\mathcal{F}_{\phi}$ except for the small region near $\phi=0$.
Common-mode noise can be a major practical limitation during the signal-acquisition (interrogation) part of a differential sensing protocol: while one wants to estimate a small phase difference between the two ensembles, the average phase of the two ensembles could be completely random and uncontrolled. In AMO sensing platforms, common-mode noise can arise from laser-phase fluctuations in optical clocks [52, 53, 54] and vibrations of reference platforms in atom interferometers [55, 56, 57]. A standard method to mitigate this noise, without assuming it to be small, is the so-called “ellipse fitting” method [58], which is originally proposed for sensing with unentangled atoms. Recent work has sought to generalize this technique by entangling each spin ensemble separately [59]. An alternative is to use an entangled state between two spin ensembles that is intrinsically insensitive to common-mode noise. As discussed in Ref. 14, the $|J=0,m=0\rangle$ state (equivalent to our $\Omega\rightarrow\infty$ steady state) is such a state, as it is invariant under global rotations. Hence, in principle its HL scaling for differential sensing is robust to such noise. However, based on Sec. III.2, the $|J=0,m=0\rangle$ state has a significant practical drawback: the lack of any net spin polarization means that metrology must employ complicated non-linear measurements.
Here, we demonstrate that our generalized class of states provide a means for combining the common-mode noise robustness of the $|J=0,m=0\rangle$ state while still allowing for a simple, linear measurement. The trick is to use our steady state $|\psi_{\rm ss}^{(2)}\rangle$ at finite values of $\Omega$, so that it retains a net spin length. The loss of rotational symmetry can be compensated by ellipse fitting, such that we can still mitigate common-phase noise and maintain the same HL scaling while measuring one-body observables. Our approach circumvents the complications of measuring two-body observables in Ref. 14.
@@
Here $\lambda_{k}$ and $|k\rangle$ are the eigenvalues and eigenvectors of $\hat{\rho}_{\rm c}$. Comparing Eq. (10) with Eq. (16), we reach a surprising conclusion: the QFI for the differential phase $\phi$ remains unchanged for any probability distribution $P(\theta)$ describing the common-mode phase noise. This result can be understood intuitively by considering the generalized Ramsey measurement scheme of the previous subsection, characterized by the two-mode spin-squeezing parameter in Eq. (12). The robustness against common phase noise is the simple consequence of the fact that $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}=\mathrm{Var}(\hat{S}^{x}_{1}+\hat{S}^{x}_{2})_{\rm ss}$. Hence, while random common phase rotations mix these collective variances, this does not alter the variance that limits the generalized Ramsey measurement. Further, as discussed in Ref. 59, this equal-variance condition corresponds to the optimal situation for extracting a differential phase $\phi$ using ellipse fitting.
@@
where $p_{j}=m_{j}/S+1/2$ with $j=1,2$. We then integrate the common phase $\theta$ over the random distribution, $P(p_{1},p_{2}|\phi)=\int d\theta P(\theta)P(p_{1},p_{2}|\phi,\theta)$, which is shown in Fig. 4(b). The CFI is thus given by
@@
In Fig. 4(c), we numerically calculate the CFI $F_{\phi}$ for $S=100$. We find that for both $\Delta/\Omega=1/\sqrt{S}$ and $\Delta/\Omega=1/S$, the CFI $F_{\phi}$ is roughly approaching the QFI $\mathcal{F}_{\phi}$ except for the small region near $\phi=0$. The best sensitivity is achieved near $\phi=\pi/4$, where Eq. (17) forms a circle. Note that one can of course still measure a small differential phase $\phi$ by deliberately adding a fixed differential phase bias of $\pi/4$.
This result shows another key feature of our scheme and the resulting steady-state solution: even with large common-phase noise, it is possible to reach Heisenberg scaling using simple Ramsey-style measurements and the ellipse-fitting method. It is also interesting to note that because of the inter-ensemble entanglement in our scheme, we achieve a much better scaling than what is possible by applying the ellipse-fitting technique to a tensor product of two single-ensemble squeezed states. As shown in Ref. 59, this only yields a scaling of CFI $F_{\phi}\sim S^{4/3}$, in contrast to the HL-type scaling $F_{\phi}\sim S^{2}$ we achieve.
@@ IV.1 Exact solution for pure steady states
We make the following ansatz for $\hat{U}_{l,l+1}$:
@@ V.1 Motivation and connection to sequential quantum circuits
Depending on the choice of drive detunings in the system Hamiltonian, output states from a variety of quantum circuits can be obtained as the dissipative steady states (see Eq. (26)). In this section, we focus on a non-trivial subset of them known as sequential quantum circuits (SQC), i.e., a linear-depth circuit with each layer acting only on one subregion in the system. SQC play an important role in the preparation of matrix product states (MPS) and other tensor-product states [61, 62, 63, 64], as well as connecting different gapped quantum phases [28]. We show that the class of steady states as outputs of SQC has a natural MPS description with a low bond dimension, and, in appropriate limits, one can achieve symmetry-protected topological (SPT) order.
@@
It is also worth to emphasize that viewing the dissipative steady state as the output state of a quantum circuit (see Eq. (26)) is a useful way of understanding the final steady state, but does not reflect the actual time-dynamics of the dissipative evolution (i.e., unlike Ref. 68 we are not directly encoding a unitary circuit into a sequence of dissipators). As we will see, the dissipative dynamics in our case is very different from the circuit dynamics, and in some cases the dissipative dynamics can provide an effective shortcut for reaching the output of the quantum circuit.
@@ V.2 MPS representation
Figure 7: (a) We consider an even number $L$ of spin-$S$ ensembles placed in a 1D spin chain, and set their detunings as $\delta_{1}=\Delta_{e}/2$, $\delta_{L}=-\Delta_{e}/2$, $\delta_{2k}=\Delta_{b}/2$, $\delta_{2k+1}=-\Delta_{b}/2$, with $k=1,2,\cdots,L/2-1$. (b) The steady-state solution is the same as the output state of a “staircase” quantum circuit: Starting from the initial dimerized state (two blue boxes linked by a bond), we apply quantum gates $\hat{U}(-\Delta_{e}/2,\Delta_{b}/2,\chi)$ (purple boxes) and $\hat{U}(-\Delta_{e}/2,-\Delta_{b}/2,\chi)$ (green boxes) alternately. The red dashed line marks the translation-invariant structure for constructing a matrix product state (MPS). (c) The upper panel shows the schematic of the spin-$1$ AKLT state, which is formed by projecting every unit cell between the spin-$1/2$ dimers to total spin $1$. The bottom panel shows the optimized relaxation timescale $t_{\rm ss}$ for reaching the steady state corresponding to fixed parameters $\Gamma$ and $\Omega/\chi$. We also set $\Delta_{e}=0$ and $\Delta_{b}=\sqrt{2}\,\chi$, which make the steady state equal to the spin-$1$ AKLT state in the large-Rabi-drive limit. The inset shows the fidelity between our steady-state solution $|\psi_{\rm ss}^{(L)}\rangle$ and the spin-$1$ AKLT state as a function of $\Omega/\chi$. The line colors specifying the system size in the inset also apply to the data points in the main plot.
@@
where the subscripts $\alpha,\beta=-S,\cdots,S$ are the internal bond-dimension indices. Calculation details and the general analytical form of the amplitudes $f_{J}(\Delta_{e},\Delta_{b},\chi)$ are discussed in App. D. Equation (32) shows that the MPS matrices factorize in the large-$\Omega$ limit: the matrix structure of $\mathbf{A}^{Jm}$ is only determined by Clebsch-Gordan coefficients, and the control parameters $\Delta_{e},\Delta_{b},\chi$ only enter through an overall $J$-dependent scalar prefactor. The form of Eq. (32) indicates that the MPS is invariant under a global on-site SO(3) symmetry [69]; at a physical level, this symmetry emerges in the large-$\Omega$ limit as each dimer in the input to our circuit is in the $|J=0,m=0\rangle$ state, and hence has no preferred axis. For a finite Rabi drive $\Omega$, the structure of the “staircase” quantum circuit remains the same with a general input state $\bigotimes_{k=1}^{L/2}|\psi_{\rm ss}^{(2)}\rangle_{2k-1,2k}$. In this case, it is more cumbersome to write down closed-form expressions for the MPS matrices, since they no longer exhibit the simple factorization in Eq. (32).
@@ V.3 Stabilization of spin-$1$ AKLT states
We see that $f_{0}=0$ can be achieved by tuning $\Delta_{e}=0$ and $\Delta_{b}=\pm\sqrt{2}\,\chi$, thus realizing the exact MPS matrices of the spin-$1$ AKLT state. As shown in the upper panel of Fig. 7(c), one can understand the spin-$1$ AKLT state as projecting every unit cell between the spin-$1/2$ dimers (singlets) to total spin $1$. It is worth mentioning that, although we have open boundary conditions, the AKLT state we dissipatively stabilized is unique. The calculation of boundary vectors in Eq. (31) is discussed in App. D.
@@
We start by considering the first case, i.e., $t_{\rm ss}$, with fixed system parameters. Similar to Sec. III.3, we define $t_{\rm ss}$ in the following way: For $t>t_{\rm ss}$, the fidelity with the steady state, $F(t)=\mathrm{tr}\Big(|\psi_{\rm ss}^{(L)}\rangle\langle\psi_{\rm ss}^{(L)}|\,\hat{\rho}(t)\Big)$, satisfies $1-F(t)<10^{-3}$. We consider the initial state where all the spins are pointing down, numerically compute $t_{\rm ss}$ using the quantum-jump method [47], and then minimize $t_{\rm ss}$ by varying the ratio $\chi/\Gamma$ with fixed $\Gamma$ and $\Omega/\chi$. Note that optimal $t_{\rm ss}$ can be achieved near $\chi\approx\Gamma$. Remarkably we find that, although our steady-state structure can be understood via an effective SQC, $t_{\rm ss}$ appears to be roughly constant with system size $L$ with a fixed value of $\Omega/\chi$ (see the bottom panel of Fig. 7(c)). This is in sharp contrast to the linear dependence on $L$ that would required for an SQC, where preparation time is simply proportional to circuit depth. We understand the difference between dissipative stabilization and the SQC based on Eq. (2), which contains interactions between every pair of ensembles instead of just nearest neighbors. Therefore, dissipative stabilization will relax the system to the steady state as a whole instead of sequential relaxation. Note that the nearly constant scaling of $t_{\rm ss}$ is also reported for dissipative stabilization of a different class of states in a related driven-dissipative setup [71]. We further explore the relaxation time scale $t_{\rm ss}$ as a function of $\Omega/\chi$ (see the bottom panel of Fig. 7(c)) and find that $t_{\rm ss}$ grows like $(\Omega/\chi)^{2}$.
@@
We next discuss the time scale $t_{\rm AKLT}$ relevant for the preparation of approximate AKLT states with a fixed fidelity with the ideal AKLT state. As shown in the inset of Fig. 7(c), for a fixed value of $\Omega/\chi$, the fidelity with $|\psi_{\rm AKLT}^{(L)}\rangle$ decreases slowly as we increase the system size $L$. To maintain the same fidelity, one has to compensate by increasing $\Omega/\chi$. We further explore this tradeoff in App. D, where we find that the preparation time for a fixed fidelity target scales as $t_{\rm AKLT}\propto L$. Our scheme serves as a promising protocol for dissipative stabilization of spin-$1$ AKLT states. Compared to other dissipative stabilization schemes for such states [42, 72, 73, 74, 75], our scheme features a more favorable scaling of $t_{\rm AKLT}$ with system size, and requires only a single dissipator, which is more directly compatible with existing cavity-QED platforms. More importantly, unlike other approaches, our scheme is capable of stabilizing more than just the AKLT state. As we have stressed, our setup is easily reconfigurable (e.g., just change drive detunings), leading to a broad class of many-body entangled steady states. We will further explore the tunability in the following subsection.

### v2

@@ III.3 Impact of single-spin dissipation during state preparation
In the presence of such unwanted dissipation, the relaxation timescale $t_{\rm ss}$ for the ideal dynamics to reach the steady state $|\psi_{\rm ss}^{(2)}\rangle$ must not be too large, otherwise the steady-state spin squeezing will be strongly degraded once we introduce $\gamma$. We define $t_{\rm ss}$ such that the ideal ($\gamma=0$) evolution yields a high fidelity with the steady state for $t>t_{\rm ss}$. More specifically  $F(t)=\mathrm{tr}\Big(|\psi_{\rm ss}^{(2)}\rangle\langle\psi_{\rm ss}^{(2)}|\,\hat{\rho}(t)\Big)$ satisfies $F(t)>1-\epsilon$ for $t>t_{\rm ss}$, where $\epsilon$ is a small positive number (we choose $\epsilon=10^{-3}$). Here, $\hat{\rho}(0)$ can be any specific state easy to initialize in an experiment. One finds that there is a fundamental tradeoff between the speed of the dynamics and the amount of steady-state squeezing and entanglement. Similar to Refs. 48, 49, one can quantify this with a lower bound on the relaxation time (see App. B),
@@
In Fig. 3(b), we compute $t_{\rm ss}$ numerically using the quantum-jump method [50] and compare with $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}$. We use an initial state $|J=2S,m=-2S\rangle$, i.e., all spins pointing down, and then minimize $t_{\rm ss}$ by varying the ratio $\Omega/\Gamma$ while keeping $\Gamma$ and $\Delta/\Omega$ fixed. The optimal $t_{\rm ss}$ can be achieved when $\Omega\gtrsim S\Gamma$. We stress that the required strong Rabi drive ($\Omega\gtrsim S\Gamma$) has been realized in cavity QED experiments with a single spin ensemble (up to $10^{4}$ atoms) [38]. Based on this choice of $\Omega$, we find $\Gamma t_{\rm ss}/(2\pi)\approx 1/\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}$, which is a constant factor away from the lower bound (see Eq. (13)). As we discussed in Sec. III.2, $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}\sim S\Delta/\Omega$, we thus have $t_{\rm ss}\sim\Omega/(S\Gamma\Delta)$.
@@ III.4 Robustness against common-mode noise during interrogation
Figure 4: (a) Schematic for differential sensing in the presence of common phase noise. We first apply the quantum channel $\mathcal{E}_{12}$ (see Eq. (3)) to stabilize the steady state $|\psi_{\rm ss}^{(2)}\rangle$. The system then evolves under the differential phase $\phi$. We next apply a global $\pi/2$ pulse (random common phase $\theta$ due to laser noise) and then perform projective measurement of each ensemble in the $z$ basis to obtain the excitation fractions $p_{1}$ and $p_{2}$. (b) Classical Fisher information $F_{\phi}$ for ellipse fitting (solid lines). The dashed lines are the corresponding quantum Fisher information $\mathcal{F}_{\phi}$. $F_{\phi}$ is roughly approaching $\mathcal{F}_{\phi}$ except for the small region near $\phi=0$. The left and right insets are color-scale plots of the probability distribution for excitation fractions $(p_{1},p_{2})$. The points $(p_{1},p_{2})$ with peak probability lie on an ellipse whose form depends on $\phi$.
Common-mode noise can be a major practical limitation during the signal-acquisition (interrogation) part of a differential sensing protocol: while one wants to estimate a small phase difference between the two ensembles, the average phase of the two ensembles could be completely random and uncontrolled. In AMO sensing platforms, common-mode noise can arise from laser-phase fluctuations in optical clocks [55, 56, 57] and vibrations of reference platforms in atom interferometers [58, 59, 60]. A standard method to mitigate this noise, without assuming it to be small, is the so-called “ellipse fitting” method [61], which is originally proposed for sensing with unentangled atoms. Recent work has sought to generalize this technique by entangling each spin ensemble separately [28]. An alternative is to use an entangled state between two spin ensembles that is intrinsically insensitive to common-mode noise. As discussed in Ref. 16, the $|J=0,m=0\rangle$ state (equivalent to our $\Omega\rightarrow\infty$ steady state) is such a state, as it is invariant under global rotations. Hence, in principle its HL scaling for differential sensing is robust to such noise. However, based on Sec. III.2, the $|J=0,m=0\rangle$ state has a significant practical drawback: the lack of any net spin polarization means that metrology must employ complicated non-linear measurements.
Here, we demonstrate that our generalized class of states provide a means for combining the common-mode noise robustness of the $|J=0,m=0\rangle$ state while still allowing for a simple, linear measurement. The trick is to use our steady state $|\psi_{\rm ss}^{(2)}\rangle$ at finite values of $\Omega$, so that it retains a net spin length. The loss of rotational symmetry can be compensated by ellipse fitting, such that we can still mitigate common-phase noise and maintain the same HL scaling while measuring one-body observables. Our approach circumvents the complications of measuring two-body observables in Ref. 16.
@@
Here $\lambda_{k}$ and $|k\rangle$ are the eigenvalues and eigenvectors of $\hat{\rho}_{\rm c}$. Comparing Eq. (10) with Eq. (16), we reach a surprising conclusion: the QFI for the differential phase $\phi$ remains unchanged for any probability distribution $P(\theta)$ describing the common-mode phase noise. This result can be understood intuitively by considering the generalized Ramsey measurement scheme of the previous subsection, characterized by the two-mode spin-squeezing parameter in Eq. (12). The robustness against common phase noise is the simple consequence of the fact that $\mathrm{Var}(\hat{S}^{y}_{1}+\hat{S}^{y}_{2})_{\rm ss}=\mathrm{Var}(\hat{S}^{x}_{1}+\hat{S}^{x}_{2})_{\rm ss}$. Hence, while random common phase rotations mix these collective variances, this does not alter the variance that limits the generalized Ramsey measurement. Further, as discussed in Ref. 28, this equal-variance condition corresponds to the optimal situation for extracting a differential phase $\phi$ using ellipse fitting.
@@
where $p_{j}=m_{j}/(2S)+1/2$ with $j=1,2$. We then integrate the common phase $\theta$ over the random distribution, $P(p_{1},p_{2}|\phi)=\int d\theta P(\theta)P(p_{1},p_{2}|\phi,\theta)$, which is shown in Fig. 4(b). The CFI is thus given by
@@
We then numerically calculate the CFI $F_{\phi}$ for $S=100$. For both $\Delta/\Omega=1/\sqrt{S}$ and $\Delta/\Omega=1/S$, the CFI $F_{\phi}$ is roughly approaching the QFI $\mathcal{F}_{\phi}$ except for the small region near $\phi=0$. The best sensitivity is achieved near $\phi=\pi/4$, where Eq. (17) forms a circle. Note that one can of course still measure a small differential phase $\phi$ by deliberately adding a fixed differential phase bias of $\pi/4$.
This result shows another key feature of our scheme and the resulting steady-state solution: even with large common-phase noise, it is possible to reach Heisenberg scaling using simple Ramsey-style measurements and the ellipse-fitting method. It is also interesting to note that because of the inter-ensemble entanglement in our scheme, we achieve a much better scaling than what is possible by applying the ellipse-fitting technique to a tensor product of two single-ensemble squeezed states. As shown in Ref. 28, this only yields a scaling of CFI $F_{\phi}\sim S^{4/3}$, in contrast to the HL-type scaling $F_{\phi}\sim S^{2}$ we achieve.
@@ IV.1 Exact solution for pure steady states
We make the following SU(2)-symmetric ansatz for $\hat{U}_{l,l+1}$:
@@ V.1 Motivation and connection to sequential quantum circuits
Depending on the choice of drive detunings in the system Hamiltonian, output states from a variety of quantum circuits can be obtained as the dissipative steady states (see Eq. (26)). In this section, we focus on a non-trivial subset of them known as sequential quantum circuits (SQC), i.e., a linear-depth circuit with each layer acting only on one subregion in the system. SQC play an important role in the preparation of matrix product states (MPS) and other tensor-product states [63, 64, 65, 66], as well as connecting different gapped quantum phases [31]. We show that the class of steady states as outputs of SQC has a natural MPS description with a low bond dimension, and, in appropriate limits, one can achieve symmetry-protected topological (SPT) order [29, 30].
@@
It is also worth to emphasize that viewing the dissipative steady state as the output state of a quantum circuit (see Eq. (26)) is a useful way of understanding the final steady state, but does not reflect the actual time-dynamics of the dissipative evolution (i.e., unlike Ref. 70 we are not directly encoding a unitary circuit into a sequence of dissipators).
@@ V.2 MPS representation
where the subscripts $\alpha,\beta=-S,\cdots,S$ are the internal bond-dimension indices. Calculation details and the general analytical form of the amplitudes $f_{J}(\Delta_{e},\Delta_{b},\chi)$ are discussed in App. D.
Equation (32) shows that the MPS matrices factorize in the large-$\Omega$ limit: the matrix structure of $\mathbf{A}^{Jm}$ is only determined by Clebsch-Gordan coefficients, and the control parameters $\Delta_{e},\Delta_{b},\chi$ only enter through an overall $J$-dependent scalar prefactor. The form of Eq. (32) indicates that the MPS is invariant under a global on-site SO(3) symmetry [71]. At a physical level, this symmetry emerges in the large-$\Omega$ limit as each dimer in the input to our circuit is in the $|J=0,m=0\rangle$ state (no preferred axis for rotation), and the unitaries in the form of Eq. (22) always preserve this symmetry. For a finite Rabi drive $\Omega$, the structure of the “staircase” quantum circuit remains the same with a general input state $\bigotimes_{k=1}^{L/2}|\psi_{\rm ss}^{(2)}\rangle_{2k-1,2k}$. In this case, it is more cumbersome to write down closed-form expressions for the MPS matrices, since they no longer exhibit the simple factorization in Eq. (32).
Figure 7: (a) We consider an even number $L$ of spin-$S$ ensembles placed in a 1D spin chain, and set their detunings as $\delta_{1}=\Delta_{e}/2$, $\delta_{L}=-\Delta_{e}/2$, $\delta_{2k}=\Delta_{b}/2$, $\delta_{2k+1}=-\Delta_{b}/2$, with $k=1,2,\cdots,L/2-1$. (b) The steady-state solution is the same as the output state of a “staircase” quantum circuit: Starting from the initial dimerized state (two blue boxes linked by a bond), we apply quantum gates $\hat{U}(-\Delta_{e}/2,\Delta_{b}/2,\chi)$ (purple boxes) and $\hat{U}(-\Delta_{e}/2,-\Delta_{b}/2,\chi)$ (green boxes) alternately. The red dashed line marks the translation-invariant structure for constructing a matrix product state (MPS). (c) The upper panel shows the schematic of the spin-$1$ AKLT state, which is formed by projecting every unit cell between the spin-$1/2$ dimers to total spin $1$. The bottom panel shows the relaxation timescale $t_{\rm ss}$ for reaching the steady state with $\chi=\Gamma$. We also set $\Delta_{e}=0$ and $\Delta_{b}=\sqrt{2}\,\chi$, which make the steady state equal to the spin-$1$ AKLT state in the large-Rabi-drive limit. The inset shows the fidelity between our steady-state solution $|\psi_{\rm ss}^{(L)}\rangle$ and the spin-$1$ AKLT state as a function of $\Omega/\chi$. The line colors specifying the system size $L$ in the inset also apply to the data points in the main plot.
@@ V.3 Stabilization of spin-$1$ AKLT states
We see that $f_{0}=0$ can be achieved by tuning $\Delta_{e}=0$ and $\Delta_{b}=\pm\sqrt{2}\,\chi$, thus realizing the exact MPS matrices of the spin-$1$ AKLT state. As shown in the upper panel of Fig. 7(c), one can understand the spin-$1$ AKLT state as projecting every unit cell between the spin-$1/2$ dimers (singlets) to total spin $1$. It is worth mentioning that the AKLT state we dissipatively stabilized is unique (see App. D and F). Nevertheless, one can still access the 4 different AKLT states with open boundary conditions depending on the projective measurement outcome of the two edge qubits [73].
@@
We start by considering the first case, i.e., $t_{\rm ss}$, with fixed system parameters. Similar to Sec. III.3, we define $t_{\rm ss}$ in the following way: For $t>t_{\rm ss}$, the fidelity with the steady state, $F(t)=\mathrm{tr}\Big(|\psi_{\rm ss}^{(L)}\rangle\langle\psi_{\rm ss}^{(L)}|\,\hat{\rho}(t)\Big)$, satisfies $1-F(t)<10^{-3}$. We consider the initial state where all the spins are pointing down, numerically compute $t_{\rm ss}$ via exact evolution of Lindblad master equation. In App. B, we show that optimal $t_{\rm ss}$ can be achieved near $\chi=\Gamma$ for fixed $\Gamma$ and $\Omega/\chi$, and the deviation from $\chi=\Gamma$ reduces as we increase $L$. Therefore, we show $t_{\rm ss}$ with $\chi=\Gamma$ as an upper bound of the optimal relaxation time scale in the bottom panel of Fig. 7(c).
Remarkably, although dissipative stabilization at $\chi=\Gamma$ (unidirectional propagation) share similarities with the SQC structure of our steady state, we find that $\Gamma t_{\rm ss}/(2\pi)\approx(\Omega/\chi)^{2}g(L)$, where $g(L)$ is a function that depends only weakly on the system size $L$, as suggested by numerical calculations up to $L=10$ (see App. B for further discussions). This is in contrast to an SQC, where preparation time (circuit depth) is simply proportional to system size $L$. As we already mentioned, the SQC with only nearest-neighbor quantum gates does not reflect the actual dissipative dynamics, since the collective dissipation and interactions in our system (see Eq. (2)) allow spins to interact with one another instantaneously, regardless of their distance. Note that the weak dependence of system size for $t_{\rm ss}$ is also reported for dissipative stabilization of a different class of states in a related driven-dissipative setup [74].
@@
We next discuss the time scale $t_{\rm AKLT}$ relevant for the preparation of approximate AKLT states with a fixed fidelity with the ideal AKLT state. As shown in the inset of Fig. 7(c), for a fixed value of $\Omega/\chi$, the fidelity with $|\psi_{\rm AKLT}^{(L)}\rangle$ decreases slowly as we increase the system size $L$. To maintain the same fidelity, one has to compensate by increasing $\Omega/\chi$. We further explore this tradeoff in App. D, where we find that the preparation time for a fixed fidelity target scales as $t_{\rm AKLT}\propto Lg(L)$. Our scheme serves as a promising protocol for dissipative stabilization of spin-$1$ AKLT states. Compared to other dissipative stabilization schemes for such states [45, 75, 76, 77, 78], our scheme features a more favorable scaling of $t_{\rm AKLT}$ with system size, and requires only a single dissipator, which is more directly compatible with existing cavity-QED platforms. More importantly, unlike other approaches, our scheme is capable of stabilizing more than just the AKLT state. As we have stressed, our setup is easily reconfigurable (e.g., just change drive detunings), leading to a broad class of many-body entangled steady states. We will further explore the tunability in the following subsection.